OpenStudy (anonymous):

Use implicit differentiation to find an equation of the tangent line to the curve at the given point x^2 + 2xy - y^2 + x =2 (1,2) (hyperbola)

OpenStudy (kc_kennylau):

x^2+2xy-y^2+x=2 2x+2x\(\frac{dy}{dx}\)+2y-2y\(\frac{dy}{dx}\)+1=0